### v1

# SemHiTok: A Unified Image Tokenizer via Semantic-Guided Hierarchical Codebook for Multimodal Understanding and Generation

Abstract
We present SemHiTok, a unified image Tokenizer via Semantic-Guided Hierarchical codebook that provides consistent discrete feature representations for multimodal understanding and generation tasks. Recently, unified multimodal large models (MLLMs) for understanding and generation have sparked exploration within research community. Previous works attempt to train a unified image tokenizer by combining loss functions for semantic feature reconstruction and pixel reconstruction. However, due to the differing levels of features prioritized by multimodal understanding and generation tasks, joint training methods face significant challenges in achieving a good trade-off. SemHiTok addresses this challenge through Semantic-Guided Hierarchical codebook which builds texture sub-codebooks on pre-trained semantic codebook. This design decouples the training of semantic reconstruction and pixel reconstruction and equips the tokenizer with low-level texture feature extraction capability without degradation of high-level semantic feature extraction ability. Our experiments demonstrate that SemHiTok achieves state-of-the-art rFID score at $256\times 256$ resolution compared to other unified tokenizers, and exhibits competitive performance on multimodal understanding and generation tasks.
1 Introduction
Figure 1: The joint tokenizer design paradigm is demonstrated. (a) represents the use of a single encoder, where semantic and texture information share the same discrete feature space in a jointly optimized manner; (b) represents the use of dual encoders, but the semantic and texture information still reside in the same discrete feature space; (c) represents the design paradigm of our method, using dual encoders and dual codebooks, achieving efficient unification of discrete features through unidirectional information interaction.
In recent years, autoregressive models have achieved great success in natural language processing, and have been extended to the multimodal understanding domain, demonstrating immense potential. This triggers researchers’ interest in unified multimodal understanding and generation by employing a single autoregressive framework. To achieve a unified multimodal large model, the key challenge is designing a tokenizer suitable for both generation and understanding tasks.
However, there is a vast gap in the visual information required for these two tasks. For instance, models from the CLIP[35, 41, 62] family, commonly used in multimodal understanding tasks, tend to lose visual texture information. On the contrary, the VQGAN[57, 67] family models, often used in autoregressive generation tasks, lack the ability to extract semantic feature for multimodal understanding. It is the core to design a unified tokenizer that simultaneously extracts textural feature and semantic feature for achieving unified MLLMs.
Recent researches combine MLLMs with diffusion models[13, 19], which introduces excessive computational. This led researchers to explore unified tokenizers. Common unified tokenizer paradigms are shown in Fig.1. VILA-U[52] and Muse-VL[55] attempt to incorporate semantic reconstruction loss during VQGAN training, aiming to derive a unified tokenizer through joint training. However, due to the gap between the two features and the optimization challenges of joint training, the model struggles to achieve a good balanced extraction of semantic and texture features. Although Janus[50] introduce a dual-encoder method that separate encoders for understanding and generation tasks to address this conflict, this increases the complexity of handling mixed tasks and does not fundamentally resolve the feature conflict challenge. TokenFlow[34] decouples semantic and pixel-level feature learning through shared mapping mechanism, whereas it does not realize a unified representation in the feature space. These limitations highlight a core challenge in the field: achieving consistent representation of semantic features and texture features, and decoupling their learning
To address this challenge, a straight forward solution is to train the semantic and texture vector quantizer separately, then concatenate quantized token either along the token length or across dimension. However, this approach increases the number of tokens thereby raising computational complexity, or results in a codebook that is too large to be suitable for MLLM. Inspired by the observation that image patches with the same semantic code tend to have similar texture feature, we propose SemHiTok, a unified image tokenizer that provides consistent feature representations for multimodal understanding and generation tasks through a unique hierarchical codebook design. The crucial point is to use same discrete features for multimodal understanding and generation tasks while decoupling the learning of semantic and texture features. Through semantic-guided hierarchical codebook, we achieve a unified representation of semantic and texture features for understanding and generation task. Unlike existing approaches that directly employ a single codebook to encode different level features, SemHiTok’s semantic-guided hierarchical codebook support stage-wise training paradigm where each stage exclusively optimizes specific hierarchy level features. This novel design allows us to achieve a better trade-off between semantic and texture feature extraction.
Our framework exhibits remarkable performance on multimodal understanding and generation tasks. For multimodal understanding, we surpass other next-token-based approaches using discrete visual inputs under the LLaVA-1.5 setting. SemHiTok also achieves competitive rFID metrics at a resolution of $256\times 256$ on ImageNet. For text-to-image generation tasks, we achieve state-of-the-art results, outperforming existing next-token autoregressive generation methods. These findings indicate that SemHiTok achieves a better trade-off between semantic and texture feature extraction capabilities, and demonstrates the great potential for application in MLLMs.
2 Related Work
2.1 Image Tokenizer
Tokenization for Generation. Image tokenizers are crucial for autoregressive image generation[45, 38, 44]. VQVAE[45] learns a discrete representation using a learnable codebook in auto-encoder architectures. VQGAN[57] further improves better perceptual quality by using adversarial training[16]. advanced the existing architecture by integrating perceptual loss and discriminator loss, alongside adversarial loss, to enhance reconstruction quality. This approach yields more precise and detailed image representations, significantly improving upon previous methodologies in image generation and processing. Subsequently, ViT-VQGAN[58] and Efficient-VQGAN[5] advance the framework with the transformer design. In recent literature, researchers are turning to efficient codebook structures[39, 63, 58] and better quantization methods[61, 59] to improve generation performance and compression rates. IBQ[39] propose the Index Backpropagation Quantization codebook update method, achieving stable training of large-scale codebooks. TiTok[59] uses a self-attention mechanism to transfer image token information to latent tokens, achieving higher compression rates and improving generation performance. Although these methods efficiently retain low-level texture information, they frequently fail to capture high-level semantic information, which limits their application to multimodal understanding tasks.
Tokenization for Understanding. In multimodal large language models (MLLMs)[26, 24, 36, 29, 1, 8], researchers leverage CLIP[36] and BLIP[24] to extract visual characteristics that align with the language during its pre-training phase. Building upon, many works[29, 1, 8] have been collected and trained on high-quality datasets to achieve remarkable performance. LLaVA[29] utilizes vision encoder to align the vision inputs before LLMs. QwenVL[1] and InterVL[8] achieve better results through increased resolution, higher-quality datasets, etc. However, these text-aligned image encoders tend to focus only on semantic information and ignore texture information, which is important for the generation task.
Figure 2: Overview of SemHiTok. We incorporate dual encoders and codebooks with guidance of SPC, which avoids the joint learning of semantic and texture features. Based on the pre-trained SPC, given an input image, we first perform semantic tokenization through SPC to obtain the encoded semantic codes. The image is then fed into pixel encoder to extract continuous texture features. Specifically, during texture feature quantization, the corresponding sub-pixel codebook is selected based on the spatially aligned semantic code. Finally, the quantized texture features, together with the quantized semantic features, are fed into the pixel decoder for pixel reconstruction.
2.2 Unified Image Tokenizer
Numerous efforts have emerged to develop unified visual generation and understanding within one MLLM[46, 48, 54, 10, 14, 42, 43, 53]. There are two main lines to bridge the gap. Many workers[10, 14, 42] combine diffusion models with LLMs for image generation. DREAMLLM[10] presents a unified framework that not only provides multimodal understanding but also creates multimodal content via diffusion models. Emu2[42] trains a unified generative model using a diffusion-based decoder. These approaches inevitably increase model complexity and are not simple enough. Other workers[43, 53, 54, 48] adopt VQVAE-based encoders to convert images into discrete tokens. Chameleon[43] and EMU3[48] directly use VQGAN[57], which is optimized by pixel reconstruction as the image tokenizer, while this method increases resource consumption during the pre-training stage and degrades multimodal understanding performance. VILA-U[53] introduces a unified image tokenizer that incorporates a text-aligned branch within the VQGAN training paradigm. However, due to the gap between semantic feature and texture feature, the joint optimization approach may lead to suboptimal solutions In contrast, our proposed SemHiTok can add the ability to extract texture features without changing the discrete semantic features, and avoids the challenges brought by joint optimization.
3 Method
The main objective of SemHiTok is to establish a simple and unified image tokenizer for understanding and generation. In this model, the image is transformed into discrete tokens that contain high-level semantic information and fine-grained texture information. We start with a Semantic-Priority Codebook(SPC) training recipe that reconstructs semantic features extracted from language image pre-training model[62, 49], and point out SPC’s poor texture feature extraction capability in section 3.1. In section 3.2, we conduct a discussion on Pixel Reconstruction Enablement(PRE), and observe that the image patches corresponding to the semantic code exhibit similar texture features. Building on this observation, we introduce Semantic-Guided Hierarchical Codebook(SGHC), incorporating texture information without compromising SPC’s understanding capability. Finally, we introduce the application of SemHiTok in multimodal downstream tasks in section 1. The overview of SemHiTok is shown in Fig.2.
3.1 Semantic-Priority Codebook Training
For multimodal understanding, use text-aligned visual encoder[62, 35, 41, 47] as image tokenizer can accelerate convergence and improve performance. However, these text-aligned visual encoder typically output continuous semantic features. In this work, to achieve a unified visual tokenizer, we first train a Semantic-Priority Codebook(SPC) to quantize the continuous semantic feature following VQKD[49].
Given an image $X^{H\times W\times 3}$, the semantic encoder $\mathcal{E}_{\text{sem}}$ extract continuous semantic features:
$Z_{sem}=\mathcal{E}_{\text{sem}}(X)\in\mathbb{R}^{h\times w\times d_{sem}}$ (1)
Where $\mathcal{E}_{\text{sem}}$ is a frozen text-aligned image encoder, e.g., CLIP[35] and SigLIP[62]. Then the quantization bottleneck $\mathcal{Q}_{sem}$ transform $Z_{sem}$ into discrete feature set space $\mathcal{C}_{sem}=\{c_{1},c_{2},...,c_{K}\}\in\mathbb{R}^{K\times d_{sem}}$. The quantization process $\mathcal{Q}_{sem}(*)$ is as follows:
$Z_{q_{sem}},I_{q_{sem}}=\mathop{\arg\min}\limits_{k\in\{1,\dots,K\}}{\|Z_{sem}-\mathcal{C}_{sem}[k]\|}$ (2)
Where $I_{q_{sem}}\in\{0,1,\dots,K-1\}^{h\times w}$ is quantized index, $Z_{q_{sem}}=\mathcal{C}_{sem}[I_{q_{sem}}]$ is discrete feature indexed from $\mathcal{C}_{sem}$. After quantization, semantic decoder $\mathcal{D}_{sem}$ maps $Z_{q_{sem}}$ to raw semantic feature space $\hat{Z}_{sem}$. The $\mathcal{Q}_{sem}$ and $\mathcal{D}_{sem}$ are end-to-end trainable by minimizing Cosine loss $\mathcal{L}_{cos}=1-\cos(Z_{q_{sem}},\hat{Z}_{sem})$ and quantization loss $\mathcal{L}_{q}(Q)$.
We conduct multimodal understanding experiments using discrete features output by SPC in the LLaVA-1.5 setting as shown in Tab.4. The experiments show that SPC is slightly inferior to continuous features but still viable. However, we further conduct an experiment reconstructing the original pixel from the quantized features extracted by SPC. The reconstructed images exhibited noticeable blurriness and a significant loss of high-frequency details, as shown in Fig. 3.
Figure 3: Visualization of reconstructed by quantized semantic features. We highlight severely erroneous regions with red dashed boxes. It is obvious that SPC’s texture feature extraction capability is poor.
3.2 Pixel Reconstruction Enablement
Discussion. In section 3.1, we demonstrate that SPC lacks the ability to extract texture information. In order to pixel reconstruction enablement and avoid a reduction of understand ability, a straightforward approach is to add an extra VQGAN[57] model. SPC extracts discrete semantic information for comprehension tasks, and VQGAN extracts discrete texture information for generative tasks, thereby avoiding the suboptimal performance brought by joint training. However, the concatenation of semantic tokens and texture tokens introduces additional challenges. There are two distinct methods of concatenation: length-wise concatenation and dimension-wise concatenation. The former approach results in an increase in the number of tokens required per image, leading to higher computational, while the latter causes the whole visual codebook size to be equal to the product of the sizes of the two codebooks, which is too large to be acceptable for LLM.
Furthermore, we present the visualization results of the semantic code, as shown in Fig.4. It can be observed that image patches corresponding to the same code exhibit similar texture features. For example, the code $\mathcal{v}_{9733}$ is more likely to be assigned to the rooster comb element in the image. At the same time, the image patches corresponding to these combs exhibit similar texture features, such as folds, patterns, and shapes Based on this observation, we propose Semantic-Guided Hierarchical Codebook to model the texture feature space corresponding to each semantic code using sub-codebook.
Figure 4: Visualization of the codes in SPC. We selected some image patches corresponding to certain codes, and it can be seen that the semantic codes in SPC have clear semantics concept, and the image patches corresponding to the same code exhibit similar low-level texture features.
Semantic-Guided Hierarchical Codebook (SGHC). The SGHC consists of SPC and several sub-codebooks, where each sub-codebook corresponds to a semantic code of SPC. Specially, given the SPC’s pre-trained semantic codebook $\mathcal{C}_{sem}=\{c_{1},c_{2},...,c_{K}\}\in\mathbb{R}^{K\times d_{sem}}$, the pixel codebook $\mathcal{C}_{pix}=\{\mathcal{C}_{pix}^{1},\mathcal{C}_{pix}^{2},\dots,\mathcal{C}_{pix}^{K}\}\in\mathbb{R}^{K\times m\times d_{pix}}$, where $\mathcal{C}_{pix}^{k}\in\mathbb{R}^{m\times d_{pix}}$ is $k_{th}$ semantic code’s pixel sub-codebook, $m$ is the sub-codebook size. At first, the SPC quantizes input image $X$ to discrete semantic token $Z_{sem}$ and semantic codebook index $I_{sem}$. In parallel, pixel encoder $\mathcal{E}_{pix}$ extract continuous texture features $Z_{pix}=\mathcal{E}_{pix}(X)$. For the quantization process of the pixel codebook, the corresponding pixel sub-codebook is selected based on the quantization result of the semantic codebook. For instance, given image patch $i$, its semantic quantization codebook index $k$ and its continuous pixel feature $Z_{pix}^{i}$, SGHC selects pixel sub-codebook $\mathcal{C}_{pix}^{k}$ to quantize $Z_{pix}^{i}$. The process is as follows:
$Z_{q_{pix}}^{i},I_{q_{pix}}^{i}=\mathop{\arg\min}\limits_{j\in\{1,\dots,m\}}{\|Z_{pix}^{i}-\mathcal{C}_{pix}^{i}[j]\|}$ (3)
where $j$ is selected sub-codebook internal index. Finally, the semantic quantized tokens and pixel quantized tokens are concatenated to $Z=concate_{dim}(Z_{q_{sem}},Z_{q_{pix}})$ as the input of pixel decoder $\mathcal{D}_{pix}$ to reconstruct raw pixel image:
$\hat{X}=\mathcal{D}_{pix}(Z)$ (4)
where $\hat{X}$ is reconstructed pixel image. The $\mathcal{E}_{pix}$, $C_{pix}$ and $\mathcal{D}_{pix}$ are end-to-end trainable by minimizing reconstruction loss $L_{img}=\ell_{1}(\hat{X},X)$, codebook loss $Lc$, perceptual loss $L_{per}$ and represents adversarial loss $L_{gan}$. The reconstruction loss is formulated as:
$L_{rec}=L_{img}+\lambda_{1}Lc+\lambda_{2}L_{per}+\lambda_{3}L_{gan}$ (5)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are loss weight of each item.
Our SGHC can be regarded as the refinement of semantic discrete space to enable pixel reconstruction. We place a specific emphases on two key advantage of SGHC: (1) Non-Conflicting Extension: Our method leverages a pre-trained semantic codebook as foundation, with pixel reconstruction losses exclusively employed to optimize pixel branch modules during the PRE. This strategic approach effectively circumvents the suboptimal solutions that arise from joint optimization processes. Furthermore, SGHC’s final output is generated by concatenating semantic-quantized features with texture-quantized features, preserving the full expressive capacity of the original semantic features while integrating complementary texture information through this unified feature fusion paradigm. It’s important to emphasize that subsequent reconstruction, multimodal understanding, and generation experiments are all conducted using discrete features derived from SGHC. (2) Efficient Downstream Applications: SGHC effectively avoids two critical predicaments: token quantity inflation and codebook overexpansion. As defined before, semantic codebook size $K$ and each pixel sub-codebook $m$. Due to dimensional splicing, the complete codebook expands to $K\times m$, where $m$ is much more smaller than $K$. In our experimental default settings, we extend complete codebook to a size comparable to existing LLM’s vocabulary size, e.g., the size of Qwen2’ vocabulary [56, 47] is $150$k. Leveraging these strengths, our framework demonstrates robust performance across multimodal downstream tasks, validating the effectiveness of the proposed architectural design.
3.3 Multimodal Downstream Tasks
Figure 5: Framework of multimodal downstream tasks.
[…]
Pixel VQGAN[57] 256 256 16,384 4.98
LlamaGen[40] 256 256 16,384 2.19
RQVAE[20] 256 256& 16,384 3.20
RQVAE[20] 256 1024& 16,384 1.30
VQGAN-LC[67] 256 256 16,384 3.01
VQGAN-LC[67] 256 256 100,000 2.62
IBQ[39] 256 256 16,384 1.37
IBQ[39] 256 256 262,144 1.00
Uni VILA-U[52] 256 1024& 16,384 1.80
SDE[55] 256 256 32,768 2.26
TokenFlow[34] 256 680 32,768 1.37
SemHiTok(ours)* 256 256 196,608 1.10
SemHiTok(ours) 256 256 196,608 1.24
Table 1: Comparison of reconstruction quality on the ImageNet-50k validation set. *: model is only trained on ImageNet. &: quantizer uses residual quantization (RQ), where the total Num are multiplied by RQ depth.
The framework diagram for multimodal downstream tasks is shown in Fig.5. We use unified discrete visual features for multimodal understanding and generation tasks Multimodal Understanding. SemHiTok acts as a visual VQ tokenizer, providing the LLM with quantized features containing semantic information. Following the LLaVA-1.5 paradigm, the image is passed through SemHiTok to obtain quantized visual features, which are then input to the LLM after a simple MLP layer. In order to achieve feature consistency, quantized semantic and texture features are spliced as input to LLM, and our experiments shown that there is negligible performance impact in comprehension tasks compared to direct utilization of quantized semantic features, confirming the viability of our feature integration strategy.
Visual Token Method # Params Res. SEEDB POPE GQA MMMU MMB MME MME-P
Continuous InstructBLIP [28] Vicuna-13B 224 58.8 78.9 49.5 50.7 36.0 1212.8
MiniGPT-4 [66] Vicuna-13B 224 1158.7 866.6
BLIP-2 [25] Vicuna-13B 224 46.4 42.5 1293.8
ShareGPT4V [7] Vicuna-7B 336 69.7 63.3 60.4 68.8 1943.8 1567.4
NExT-GPT [51] Vicuna-7B 224 57.5 58.0
Qwen-VL-Chat [2] Qwen-7B 448 57.7 57.5 1848.3 1487.5
Janus [50] DeepSeek-LLM-1.3B 384 63.7 87.0 59.1 30.5 69.4 1338.0
LLaVA-1.5 [30] Vicuna-13B 336 68.1 85.9 63.3 36.4 67.7 1826.7 1531.3
Discrete LWM [31] LLaMA-2-7B 256 75.2 44.8
SEED-LLaMA [22] LLaMA-2-13B 224 53.7
Show-o [54] Phi-1.5-1.3B 256 80.0 58.0 26.7 1097.2
EMU3 [48] 8B from scratch 512 68.2 85.2 60.3 31.6 58.5 1509.9 1243.8
MUSE-VL[55] Qwen2.5-7B 256 70.0 42.3 72.1 1480.4
VILA-U [52] LLaMA-2-7B 256 56.3 83.9 58.3 1336.2
VILA-U [52] LLaMA-2-7B 384 78.7 86.8 1401.8
TokenFLow-L [52] Vicuna-13B 256 62.6 85.0 60.3 34.4 60.3 1622.9 1365.4
TokenFLow-XL [52] Vicuna-13B 384 59.0 86.8 62.7 38.7 68.9 1590.4 1401.8
SemHiTok(ours) Vicuna-7B 256 58.7 83.2 58.8 34 54.7 1590.4 1307.4
SemHiTok(ours) Vicuna-7B 384 62.4 84.2 61.0 35.2 60.3 1724.2 1400.6
Table 2: Evaluation on multimodal understanding benchmarks. We include approaches with continuous visual inputs versus discrete visual inputs. SemHiTok achieved competitive performance under the LLaVA-v1.5 setting.
Visual Generation. We deploy the visual generation model on pre-trained LLM. For encoding text, we directly use LLM’s BPE tokenizer to convert text into discrete token sequences. We add special visual tokens to the original vocabulary and adjust the head layer accordingly. Follow the autoregressive image generation training paradigm, we leverage SemHitok to convert images into discrete token sequences, and splice them with text token sequences for input to LLM, using the cross-entropy function to compute the loss of visual tokens. Regarding the discrete coding of vision, for example, for the $j_{th}$ code in the $i_{th}$ pixel sub-codebook, the discrete code index in the completed codebook is $h=i\times m+j$. To enable classifier-free guidance[17], we randomly replace the text condition with probability of 0.1 to the un-condition text during training.
4 Experiments
4.1 Experimental Setup
Datasets. The training of SemHiTok is conducted in two stages. During SPC training, we train the semantic tokenizer for one epoch on 50M subset of COYO-700M[4]. For PRE stage, we first trained the model on ImageNet[9], and then fine-tune on 50M COYO-700M subset to improve its generalization, following LlamaGen[40]. To further improve the reconstruction and generation performance, we enlarge the size of the pixel decoder and fine-tune it on the same data. For multimodel understanding, we train VLM on LLaVA-Pretrain558K and LLaVA-v1.5-mix-665K. To evaluate the performance of the VLM, we test the model on vision-language benchmarks such as SEEDB: SEED-Image[21], POPE[27], MMMU[60], MMB[32], MME-P: MME-Perception[12] and MME[12]. We run the evaluation based on lmms-eval[64]. For visual generation with SemHiTok, we use 15M in-house text-image pairs,and evaluate on GenEval[15] and MJHQ30K[23].
Implement Details. For SPC, we employ SigLIP[62] as the semantic encoder and update semantic codebook using EMA[18] algorithm to enhance stability. We use SigLIP-large-p16 and SigLIP-so400m-p14 as semantic encoder for $256\times 256$ and $384\times 384$ resolution, respectively. We set semantic codebook size is $16384$. The semantic decoder comprises three self-attention layers[11]. In the pixel branch, we first use ViT-Base as the pixel encoder and decoder, then increase the pixel decoder to a ViT-Large in the fine-tuning enhanced decoder stage. The default pixel sub-codebook size is $12$, so the whole codebook size is $196608$. For multimodal downstream tasks, we use Vicuna-v1.5-7B[29] and Qwen2.5-3B[1] as the language model for multimodal understanding tasks and generation tasks, respectively. During image generation inference, we apply classifier-free guidance[17], with a scale factor of $2$.
4.2 Unified Image Tokenizer
We present reconstruction performance of SemHiTok on ImageNet-50k validation set at $256\times 256$ resolutions and 16×16$\times$16 × compression ratio in Tab.1. The * indicates that model is only trained on ImageNet, which is the same as the expert pixel reconstruction. In an ImageNet-only training setting, SemHiTok achieves 1.10 rFID, which outperforms the most of pixel-expert models, such as LlamaGen and RQVAE. Notably, compared to IBQ which has close codebook size to ours, SemHiTok is only higher by 0.1.
Compared with the unified tokenizer, SemHiTok achieves 1.24 rFID which is state-of-the-art performance. These results demonstrate the effectiveness of our SGHC design scheme. In the subsequent multimodal understanding experiments and visual generation experiments, we use the model in the last column as the image tokenizer.
Model Text Pretrain Res. GenEval MJHQ30K
Overall $\uparrow$ gFID $\downarrow$
SD v1.5 [37] CLIP ViT-L/14 512 0.43
SD v2.1 [37] CLIP ViT-H/14 768 0.50
SDXL [33] CLIP ViT-bigG 1024 0.55 9.55
PixArt-alpha [6] Flan-T5-XXL 512 0.48 6.14
DALL-E 3 [3] Flan-T5-XXL 1024 0.67
Show-o [54] Phi-1.5 256 0.53 15.18
Transfusion [65] 256 0.63
LWM [31] 256 17.77
LlamaGen [40] Flan-T5-XL 512 0.32
EMU3 [48] 512 0.54
Janus [48] 384 0.61 10.10
VILA-U [52] 256 12.81
SemHiTok(ours) 256 0.66 11.0
Table 3: Comparison of generation quality on GenEval [15] and MJHQ30K [23].
4.3 Multimodal Understanding
We present the performance of VLMs using SemHiTok’s encoder on vision-language benchmarks in Tab. 2. VLMs are classified into two categories based on whether the input visual token is discrete: continuous visual tokens and discrete visual tokens. Compared to VLMs based on continuous tokens, LLMs utilizing discrete tokens generally exhibit lower performance. Among LLMs with discrete tokens, SemHiTok achieves competitive performance despite being trained on a smaller dataset and built upon a old-fashioned base model. To be specifically, compared to LWM, which employs a traditional VQVAE and uses same size LLM as ours, SemHiTok significantly outperforms it across all metrics. Additionally, when compared to VILA-U, which also adopts a unified tokenizer, SemHiTok achieves comparable performance. Our experiments demonstrate that SemHiTok effectively achieves a satisfactory trade-off between semantic and texture information.
4.4 Text-to-Image Generation
Figure 6: Presentation of the generated results at a resolution of $256\times 256$ on MJHQ30K. Our generated images show good consistency with the prompts in terms of objects and style
The results of SemHiTok on the text-to-image generation task are shown in Tab. 3. On GenEval, SemHiTok surpasses all autoregressive generation models and most diffusion models, such as SDXL, and is very close to the best-performing DALL-E. This indicates that our generation model is able to align with the prompt at the instance level. We attribute this to the powerful text feature extraction capabilities of the Qwen2.5 LLM, as well as the SGHC, which contains semantic information and can efficiently align the text feature space, ensuring accuracy in object generation. On the MJHQ30K, our model achieved 11.0 gFID and also demonstrated competitive performance, exceeds show-o and VILA-U. In Fig. 6, we present some generated images from MJHQ30K. Our model also demonstrates consistency with the text on MJHQ30K, even with long prompts. Owing to the powerful reconstruction capabilities of SemHiTok, the generated images are rich in detail and highly realistic, such as the fur of human and animal, the texture of leaves, and so on.
4.5 Ablation
Impact of Key Design. In Tab.4, we validate the impact of our key design choices in SemHiTok: initialization for the semantic encoder, incorporation of texture features, semantic-guided hierarchical codebook(SGHC). The baseline experimental setup involves training SPC with a randomly initialized semantic encoder. Incorporation of texture features refers to the joint training of texture information based on SPC, following the setup of MUSE-VL[55].
Our experiments indicate that initialization for the semantic encode with a teacher model significantly enhances the model’s understanding performance. Meanwhile, the semantic tokenizer that only considers semantic information fails to effectively capture the texture information for reconstruction. Furthermore, although a simple joint training approach improves rFID reduced by 0.63, it leads to a decline in understanding performance which is non-negligible. Finally, with the introduction of the SGHC design, not only is reconstruction performance improved, but understanding ability is also preserved without degradation. These experiments highlight the core advantage of the SGHC, which enables the incorporation of texture feature extraction without compromising the understanding ability of the semantic tokenizer.
SemEnc Init. Texture Incorp. SGHC. rFID $\downarrow$ GQA $\uparrow$ POPE $\uparrow$ SEEDB $\uparrow$
3.10 56.7 80.4 54.1
[…]
Table 4: Impact of key design choices on reconstruction quality and multimodal understanding benchmarks. The bolded parts represent the best results.
Impact of Fusion Type and Sub-Codebook Size. For efficiency, we only conduct training and evaluation on ImageNet. We experiment to investigate the impact of the fusion type of semantic and pixel discrete features, as well as the sub-codebook size, on reconstruction performance as shown in Tab. 5. The gray bar represents the default setting in our experiments. For fusion type, experiments show that concatenation along the dimension yields the best performance. In contrast, the reconstruction performance without semantic discrete token is the worst, indicating that semantic information can be used as supplementary information to improve the reconstruction. Due to semantic discrete tokens correspond to texture discrete tokens in 2D grid space, the concatenation fusion type in the length direction is not appropriate. For sub-codebook size, increasing the size can improve the model’s reconstruction performance, but it exhibits marginal utility. In addition, the codebook usage significantly decreases when the sub-codebook size is set to 16, which indicates that too large a sub-codebook size is not cost-effective.
[…]
w/o sem 12 1.99 98.4%
Len 12 1.45 97.1%
Dim 8 1.42 96.4%
12 1.26 93.7%
16 1.19 79.3%
Table 5: Impact of fusion type and sub-codebook size. w/o sem: Not use semantic discrete token. Len: concatenate in length direction. Dim: concatenate in dimension direction
5 Conclusion
In this work, we introduce SemHiTok, a unified image tokenizer that implements better trade-off between high-level and low-level semantic information. SemHiTok innovatively utilizes semanitc-guied hierarchical codebook to realize the reconstruction capability of texture features without affecting the understanding performance of the original semantic codebook, and achieves SOTA at ImageNet $256\times 256$ resolution. We conducted extensive experiments on multimodal understanding tasks and generation tasks, and SemHiTok achieves competitive performance with a small amount of data. These results verify the powerful unified multimodal capability of SemHiTok, and provide an effective solution for understanding and generating integrated multimodal large models.
Limitation In this work, we primarily focus on pixel reconstruction enablement without any degeneration in understanding performance based on the existing semantic codebook, without extensively exploring the design of the semantic codebook. In addition, we mainly conducted experiments on multimodal downstream tasks at a resolution of $256\times 256$, and further exploration is needed to assess performance variations at higher resolutions.
Future Work In the future, we can explore more efficient training methods and advance codebook structure for semantic tokenizer to bridge the gap between discrete and continuous features in multimodal tasks.